OpenStudy (shylovesyou11):

I just need some help on these problems.. Please? 1) What are the x- and y-intercepts of the line? 2x + 3y = –12 A) x-intercept –6, y-intercept –4 B) x-intercept –6, y-intercept 4 --My answer C)x-intercept –4, y-intercept 6 D)x-intercept 4, y-intercept –6 2) What is the slope of the line that passes through the points (6, 4) and (3, 8)? A) 4\3 B)-3\4 C) 3\4 D) -4\3 * I am not sure what this answer would be..

OpenStudy (anonymous):

x=-6 y=-4

OpenStudy (anonymous):

8-4=4 3-6=-3 =-4/3 D

OpenStudy (shylovesyou11):

Ahhhh Thank you again! =)

OpenStudy (anonymous):

welcomes :)